A hollow disk is rotating about its center at 5 rad/s. Friction is acting on the disk at the axis of rotation producing an angular acceleration of 1 rad/s.

How many revolutions will the disk rotate before coming to a stop?
Answer #1

Solution) wo = 5 rad/s

Angular acceleration = 1 rad/s^2

w = 0

Theeta = ? ( number of revolutions )

w = wo - (angular acceleration)(t)

0 = 5 - 1(t)

t = 5 s

Theeta = (wo)(t) - (1/2)(angular acceleration) (t^2)

Theeta = 5×5 - (1/2)(1)(5^2)

Theeta = 12.5 rad

1 rev = 2(pi) rad

1 rad = 1/(2(pi)) rev

Theeta = (12.5)/(2(pi))

Theeta = 1.98 revolutions